Find the critical t-value(s) for a two independent samples t-test given:

α = 0.05

n1 = 12

n2 = 11

two-tailed test
Answer #1

Degree of freedom=n1+n2-2

=12+11-2=21

g= 0,05

Critical Value(t_{\alpha/2})=(-2.08, 2.08)

(use excel function:=TINV(0.05,21))

Ans:Critical Value(t_{\alpha/2})=(-2.08, 2.08)
Answer #2

Use Table C.2 to find the critical value for a two-tailed independent samples t-test where α=.05 and the degrees of freedom are equal to 15. Report the value exactly as it appears in the table (i.e., X.XXX).